OpenStudy (anonymous):

A sphere is charged with electrons to 7.18×10−6C. The charge of an electron is 1.6×10−19C. How many electrons are there on the sphere?

OpenStudy (ranga):

Number of electrons = \(\Large \frac{7.18 \times 10^-6}{1.6 \times 10^{-19}} = ? \)

OpenStudy (kainui):

If I say I have a weighing scale that reads 10 pounds with a bunch of soda cans on it and I say a soda can weighs 2 pounds then that means I must have 5 cans of soda on it right? It's really the exact same problem.

OpenStudy (anonymous):

4.4875?

OpenStudy (ranga):

\(\Large 4.4875 \times 10^{-6 + 19} = 4.4875 \times 10^{13} ~ \text{electrons.}\)

OpenStudy (anonymous):

44875000000000?

OpenStudy (ranga):

Just leave it in scientific notation as: \(\Large 4.4875 \times 10^{13} ~ \text{electrons.}\)